OpenStudy (allen27):

Write an equation of a line parallel to y = x - 6 in slope-intercept form that passes through the point (-1, 5). (5 points) y= x - 1 y= x + 5.25 y= -4x + 5 y= -4x - 1.25

OpenStudy (allen27):

@mayankdevnani

TheSmartOne (thesmartone):

I think you posted the same question earlier.

OpenStudy (allen27):

i got no anwser tho

TheSmartOne (thesmartone):

I see, well parallel lines have the same slope so what is the slope in the equation y = x - 6? It's in the form y = mx + b, where m is the slope

TheSmartOne (thesmartone):

another hint: x is the same thing as 1x